# Scientific Literature Hierarchy Construction and Organization

### v1

@@ Abstract
Scientific knowledge is growing rapidly, making it challenging to track progress and high-level conceptual links across broad disciplines. While existing tools like citation networks and search engines make it easy to access a few related papers, they fundamentally lack the flexible abstraction needed to represent the density of activity in various scientific subfields.
We motivate Science Hierarchography, the goal of organizing scientific literature into a high-quality hierarchical structure that allows for the categorization of scientific work across varying levels of abstraction, from very broad fields to very specific studies. Such a representation can provide insights into which fields are well-explored and which are under-explored.
To achieve the goals of Science Hierarchography, we develop a range of algorithms. Our primary approach combines fast embedding-based clustering with LLM-based prompting to balance the computational efficiency of embedding methods with the semantic precision offered by LLM prompting. We demonstrate that this approach offers the best trade-off between quality and speed compared to methods that heavily rely on LLM prompting, such as iterative tree construction with LLMs. To better reflect the interdisciplinary and multifaceted nature of research papers, our hierarchy captures multiple dimensions of categorization beyond simple topic labels. We evaluate the utility of our framework by assessing how effectively an LLM-based agent can locate target papers using the hierarchy. Results show that this structured approach enhances interpretability, supports trend discovery, and offers an alternative pathway for exploring scientific literature beyond traditional search methods.111 Code, data and demo: https://github.com/JHU-CLSP/science-hierarchography
@@ 1 Introduction
The publication of scholarly work has been increasing significantly in recent years (Ware & Mabe, 2015). Despite the overall growth, this expansion is not uniformly distributed across all scientific fields. In fact, certain areas likely receive more attention than others, presenting both opportunities and challenges (Hope et al., 2023). Therefore, a natural question is:
do we have a mechanism to understand the distribution of scientific efforts across different sub-areas?
Answering this question is crucial, as academic institutions can identify emerging or underexplored research areas, and policymakers can make more informed decisions about funding allocation based on the actual distribution of scholarly work, potentially directing more resources to important but under-explored areas.
Among existing solutions, conventional tools for discovering scientific literature, search engines like Google Scholar primarily function as retrieval engines. They are designed to find the few, most relevant literature matching a specific query rather than providing a comprehensive, structured overview of the entire scientific landscape. Similarly, modern LLM-based chatbots can recite related works if they were observed during pre-training, but again it’s not easy to glean a bird’s-eye view of scientific progress.
@@
Figure 1: An example of  Science Hierarchography which illustrates the organization of scholarly work. Starting from broad research fields at the top levels, narrowing down through sub-clusters, and ultimately reaching specific papers within a particular research domain at the bottom level. Crucially, this must be inferred at scale and automatically.
Understanding the distribution properties of science fundamentally requires a mechanism to abstract over scientific concepts (such as open science problems and techniques) and connect these abstractions to various, more concrete instances. At one end of the spectrum, we have broad scientific communities (like physics or artificial intelligence), and at the most specific level, we have papers published in these areas. Between these two endpoints lie intermediate concepts that group together sub-areas. Essentially, we need a data structure that encompasses both these endpoints of abstraction and various in-between abstractions.
@@
We propose the construction of large-scale hierarchical representations of scientific literature, which we refer to as Science Hierarchography. A well-designed hierarchy offers several benefits: it provides a macro-level view of scientific progress, showing how research efforts are distributed across different methodologies and application areas. This insight helps researchers identify emerging trends and gaps, while aiding policymakers in more strategic resource allocation. Additionally, such a hierarchy offers an alternative way to explore literature, complementing traditional retrieval systems by enabling users to navigate science through conceptual hierarchies rather than isolated keyword searches.
@@
How should one represent scholarly works? A key challenge in building a hierarchical structure for scientific literature is defining the unit of representation for individual nodes. Research papers often span multiple topics (for example, reinforcement learning applied to medical imaging” or deep learning for oceanography”). We devise a prompting strategy to break down papers along multiple key contribution types (such as the problems they tackle and the techniques they address), capturing their multifaceted nature. For a fixed contribution type from science papers, we construct their hierarchical representation. Our approach ensures that papers are structured under meaningful and coherent categories (§3.2).
@@
What are construction strategies that allow for both scalability and quality? To tackle this goal, we propose Scychic (pronounced “psychic”), a novel approach to constructing a high-quality hierarchical structure for organizing scientific literature. Scychic combines fast embedding-based clustering with LLM-based prompting to balance the computational efficiency of embedding methods with the semantic precision of LLM prompting (§4.1).
How can we assess the quality of a given science hierarchy? There is no definitive ground truth for a scientific hierarchy, as it naturally evolves over time—today’s scientific hierarchy differs from that of last year. We approach our evaluation through utilization, focusing on whether an information seeker (human or AI agent) can efficiently locate specific information (such as details related to child nodes) by navigating the hierarchy from its root. Our evaluation is based on the premise that an effective hierarchy facilitates rapid information discovery, even though its applications extend beyond mere discovery (§5.2).
@@
What did we empirically observe? We demonstrate that this approach offers the best trade-off between quality and speed compared to methods that heavily rely on LLM prompting, such as iterative tree construction or iterative pruning algorithms with LLMs. Through extensive experiments, we show that Scychic, our method, can construct higher-quality hierarchies compared to a wide range of baseline approaches (§5.4). Meanwhile, validation of our approach on a $10K$ paper dataset demonstrates robust accuracy, confirming its scalability for large-scale applications.
Contributions: (1) We propose the goal of constructing large-scale abstract hierarchies for scientific literature, which can provide essential insights into the distribution of scholarly work across various fields. (2) We establish an evaluation framework that frames it as utilization—effective information discovery by traversal of the hierarchy. (3) We present Scychic, a novel approach that integrates fast embedding-based clustering with LLM-based prompting to develop a high-quality hierarchical structure, tailored to capture different dimensions of academic work. Our extensive experiments demonstrate that our approach constructs superior hierarchies compared to existing baseline methods, ultimately paving the way for a more organized and insightful understanding of scientific progress.
@@ 2 Related Work
Approaches to hierarchy induction: Hierarchy (taxonomy) induction has evolved significantly from early pattern-based methods to sophisticated, LLM-augmented frameworks. The pioneering work of Hearst (1992) initiated the field by leveraging hand-crafted hyponym patterns to bootstrap is-a relation extraction. Subsequent studies advanced these ideas by combining statistical techniques with large-scale information extraction to automatically identify hypernym-hyponym patterns (Pantel & Pennacchiotti, 2006; Yang & Callan, 2009; Girju et al., 2006). More recent efforts have incorporated LLMs and iterative prompting to refine and complete taxonomies; for example, Wan et al. (2024); Zeng et al. (2024a); Chen et al. (2023); Zeng et al. (2024b) utilized zero-shot/few-shot reasoning and ensemble-based ranking for hierarchy construction. Complementary approaches include autonomous creation of fine-grained taxonomies without fixed vocabularies (Gunn et al., 2024), self-supervised expansion in low-resource domains (Mishra et al., 2024), and graph-based methods that integrate metadata and citation information (Cong et al., 2024; Sas & Capiluppi, 2024; Shen et al., 2024). Additional works further demonstrate the effectiveness of optimization and in-context learning (Hu et al., 2024b; Shi et al., 2024; Xu et al., 2025; Jain & Espinosa Anke, 2022; Chen et al., 2021). Our work differs from existing literature in problem scope, scale, and algorithmic solutions. Our aim is to progress toward scaling taxonomy induction for the domain of scholarly works, which is more challenging than prior domains in taxonomy induction (e.g., entity names) due to the complexity of scientific works, their size and evolving nature. The closest work to ours is by Oarga et al. (2024), but our broader scope requires different algorithmic approaches and lacks ground truth labels.
@@
Structured representation of science: Scientific knowledge is expanding at an unprecedented pace (Teufel et al., 1999; Pertsas & Constantopoulos, 2017; Constantin et al., 2016; Fisas et al., 2016; Liakata et al., 2010), prompting the development of various frameworks for constructing structured representations and knowledge graphs (Fathalla et al., 2017; Jaradeh et al., 2019; Oelen et al., 2020; Vogt et al., 2020; Soldatova & King, 2006). For instance, Pham et al. (2024) develop a prompt-based method for generating interpretable topics with natural language labels, while Cui et al. (2024) and Marchenko & Dvoichenkov (2024) propose iterative frameworks that integrate object properties and graph mining techniques to refine taxonomies. Similarly, Zimmermann et al. (2024) combine curated otologies with data-driven topic maps to offer an intuitive, multi-topic navigation system. Building on and extending these works, our work focuses on constructing a high-quality hierarchical structure for the scientific literature domain.
@@ 3 Science Hierarchography: Toward Hierarchy of Scholarly Work
We begin with a formal definition of the problem (§3.1), introduce our evaluation (§5.2), and discuss the design choices on representing scientific content in our hierarchy (Section 3.2).
@@ 3.1 Formal Problem Statement
Formally, this goal is defined as inferring a graph that encodes relationships between nodes (atomic concepts representing scholarly ideas or goals). This can be formally defined as a tuple $T\vcentcolon=(N,E)$ where:
  • $N$ is the set of all concepts (nodes) present in the hierarchy: $N=\{n_{1},n_{2},\ldots,n_{N}\}$.
@@
[…]
@@
The edge (relations connecting two nodes) have types that is either “isA” relationships or “instanceOf.” The ”isA” relationship defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., ”RLHF isA RL” means ”RLHF” is a type of ”RL”). The ”instanceOf” relationship indicates that a specific instance belongs to a concept category (e.g., ”Ouyang et al. (2022) instanceOf RLHF” means ”Ouyang et al. (2022)” is a specific instance of the class ”RLHF”).
3.2 Decomposing Each Science Paper into Atomic Contributions
A key question is how to represent scholarly work content within hierarchy nodes. Science documents are rich in ideas, covering broad goals, specific problems, and techniques. This necessitates extracting structured representations of papers (D’Souza & Auer, 2020) to disentangle their different focuses. This also addresses length challenges: Most papers are 4 to 10 pages, spanning 5K to $10K$ tokens. Processing all this information for many papers would be prohibitively challenging and costly for any language model.
@@ 3.2 Decomposing Each Science Paper into Atomic Contributions
We use an LLM to preprocess each paper (title and abstract) and break them down into a pre-defined set of contributions. We consider the following types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes). Prompts for extracting contributions and examples are detailed in §C. We note that each contribution may include additional dimensions (sub-contributions). For instance, a “result” encompasses both the “outcome” and its “potential impact.”
3.3 Choosing the Effective Number of Layers in the Hierarchy
Although the effective number of layers is an empirical question, we can derive useful intuition by considering the structure of a near-balanced tree. Suppose the tree has $L$ layers and a branching factor of $b$. Such a tree will contain approximately $O(b^{L})$ nodes. If we aim to organize $C$ contributions into a near-balanced hierarchy, we must ensure that the total number of leaf and intermediate nodes grows proportionally with $C$. To achieve this, the required depth $L$ must scale logarithmically with the number of contributions, specifically as $L=O(\log_{b}C)$. In our experiments, we selected $L=3$ for a corpus of $2K$ papers and $L=4$ for $10K$ papers, both aligning with this logarithmic relationship. Extrapolating from this trend, if we were to scale up the corpus size to a range of $10^{6}$ to $10^{7}$ papers, we would expect the necessary depth to increase to approximately $L=6$ or $L=7$.
4 A Spectrum of Approaches to Tackle Science Hierarchography
@@
We present algorithms to address our proposed goal. We begin with our approach (Scychic; §4.1), its special cases (§4.2) and then discuss other baselines that heavily rely on LLMs (fLMSci; §4.3). While all these methods utilize LLMs, they vary in their reliance on them; some require many LLM calls (linear or quadratic relative to the number of papers), while others require fewer (logarithmic relative to the number of papers). Given the aim to scale to millions of papers, scalability is crucial. The objective is to find the approach that produces the highest quality hierarchy with the fewest LLM calls, which are costly in terms of latency and finances.
@@ 4.1 Scychic: Alternating Between Clustering and Summarization
7:         for each cluster from layer $l-1$ do
@@
For the first layer $L=1$, we start by dividing the entire collection of papers into $k_{1}$ clusters at the top level using their embeddings. For the second level, we process each top-level cluster independently, i.e., within each top-level cluster, we take only the papers belonging to that cluster and apply clusterer to their embeddings to divide them into subgroups.
We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$. To build layer $l$, we start by embedding the generated cluster summaries from the level below $l-1$ using embedder, similar to how we embedded the papers. We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters. This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.

### v2

@@ Abstract
Scientific knowledge is growing rapidly, making it difficult to track progress and high-level conceptual links across broad disciplines. While tools like citation networks and search engines help retrieve related papers, they lack the abstraction needed to capture the density and structure of activity across subfields.
We motivate Science Hierarchography, the goal of organizing scientific literature into a high-quality hierarchical structure that spans multiple levels of abstraction—from broad domains to specific studies. Such a representation can provide insights into which fields are well-explored and which are under-explored. To achieve this goal, we develop a hybrid approach that combines efficient embedding-based clustering with LLM-based prompting, striking a balance between scalability and semantic precision. Compared to LLM-heavy methods like iterative tree construction, our approach achieves superior quality-speed trade-offs. Our hierarchies capture different dimensions of research contributions, reflecting the interdisciplinary and multifaceted nature of modern science. We evaluate its utility by measuring how effectively an LLM-based agent can navigate the hierarchy to locate target papers. Results show that our method improves interpretability and offers an alternative pathway for exploring scientific literature beyond traditional search methods. 111 Code, data and demo are available: https://github.com/JHU-CLSP/science-hierarchography
Science Hierarchography:
Hierarchical Organization of Science Literature
Muhan Gao, Jash Shah, Weiqi Wang, Daniel Khashabi Department of Computer Science Johns Hopkins University Baltimore, MD 21218, USA {mgao38,jshah48,wwang194,danielk}@jhu.edu
@@ 1 Introduction
The pace of scientific publishing is accelerating (Ware and Mabe, 2015), but this growth is uneven across fields (Hope et al., 2023). Some areas attract dense research activity, while others remain underexplored. This raises a natural question:
How do we understand the distribution of scientific efforts across different sub-areas?
Answering this question is essential for both academic and policy stakeholders. A clearer view of how research efforts are distributed enables institutions to spot emerging or neglected areas, prioritize strategic hiring and future agendas. For policymakers, it supports more informed funding decisions, ensuring that critical but underexplored domains receive the attention and resources they deserve.
@@
Figure 1: An example of  Science Hierarchography illustrates how scholarly work can be organized hierarchically—from broad research domains at the top, through increasingly specific sub-clusters, down to individual papers at the lowest level. Critically, this structure must be inferred automatically and at scale.
Conventional tools like Google Scholar are designed as retrieval engines, optimized to return a handful of papers that match a specific query. They offer little in the way of a comprehensive or structured view of the broader scientific landscape. Similarly, while modern LLM-based assistants can surface related works (seen during pretraining or via their retrieval tools), they fall short in offering a broad, bird’s-eye perspective on scientific progress.
Addressing this challenge requires abstraction: a way to generalize over research problems and techniques and to connect broad scientific areas to specific papers via intermediate categories. At one end, we have high-level domains (e.g., physics, AI); at the other, individual papers. Between them lie a latent spectrum of subfields and methodological clusters. What’s missing is a data structure that captures all these abstraction levels.
@@
We propose building large-scale hierarchical representations of scientific literature, which we call Science Hierarchography. A well-designed hierarchy provides a macro-level view of scientific progress, revealing how research is distributed across methods and application areas. This helps researchers spot emerging trends and gaps, and supports policymakers and institutions in making more strategic resource decisions. It also offers a new way to explore the literature—complementing traditional search by allowing users to navigate science through conceptual hierarchies.
@@
How should scholarly work be represented? A central challenge in building a scientific hierarchy is defining what each node represents. Research papers often span multiple topics (e.g., reinforcement learning for medical imaging or deep learning for oceanography). To capture this complexity, we develop a prompting strategy that decomposes papers into key contribution types—such as the problems addressed and techniques used3.2). For each fixed contribution type, we construct a corresponding hierarchical structure, ensuring that papers are organized into meaningful, coherent categories.
@@
What construction strategies balance scalability and quality? To address this, we introduce Scychic (pronounced “psychic”), a new method for building high-quality hierarchical structures of scientific literature. Scychic integrates fast embedding-based clustering with LLM prompting, combining the efficiency of embeddings with the semantic precision of language models (§4.1).
How can we evaluate the quality of a scientific hierarchy? Scientific hierarchies lack a fixed ground truth—they evolve over time as research landscapes shift. We therefore adopt an evaluation-through-utilization approach, measuring whether an information seeker (human or AI) can efficiently locate specific content (e.g., child nodes) by navigating the hierarchy from the root. This evaluation hinges on the idea that a good hierarchy enables rapid information discovery, even though its utility extends well beyond search alone (§5.2).
@@
What did our empirical results show? Our approach achieves the best trade-off between quality and speed when compared to LLM-heavy methods like iterative tree construction or pruning. Extensive experiments show that Scychic consistently produces higher-quality hierarchies than a broad set of baselines (§5.4). Validation on a 10K-paper dataset further confirms its strong accuracy and scalability for large-scale use.
Contributions: (1) We introduce the goal of constructing large-scale, abstract hierarchies of scientific literature to reveal how scholarly efforts are distributed across research areas. (2) We propose a utilization-based evaluation framework that measures how effectively users can discover information by traversing the hierarchy. (3) We present Scychic, a new method that combines fast embedding-based clustering with LLM prompting to build high-quality, multidimensional hierarchies. Extensive experiments show that Scychic outperforms baseline approaches, offering a more structured and bird’s-eye view of scientific progress.
System # of Levels Node content Node granularity Assigned by Purpose Public
Web of Science One Research areas One keyword Publisher Indexing No
Scopus Two Research areas One keyword Editor Indexing Yes
arXiv Taxonomy Two Research areas One keyword Authors Indexing Yes
PubMed MeSH Multiple Medical headings One keyword Authors Indexing Yes
Microsoft Academic Graph Multiple Research areas Multiple keywords Algorithms Indexing Discontinued
Science Hierarchography (Ours) Multiple (by designer) Rich contribution descriptions Science contribution summary (many tokens) Algorithms Exploratory Analysis Yes
Table 1: Comparison of hierarchical resources for organizing scientific literature, ordered by hierarchy depth. Conventional systems are built for indexing, relying on fixed, shallow taxonomies with keyword-based nodes and human-assigned labels. In contrast, Science Hierarchography supports deeper, designer-controlled hierarchies with rich natural-language summaries, enabling more flexible and exploratory analysis of scientific work.
@@ 2 Related Work
Gierarchy induction: The field of taxonomy induction has progressed from early pattern-based techniques to modern LLM-augmented methods. Seminal work by Hearst (1992) introduced the use of hand-crafted hyponym patterns for extracting is-a relationships. Subsequent research expanded on this using statistical methods and large-scale information extraction to identify hypernym-hyponym structures (Pantel and Pennacchiotti, 2006; Yang and Callan, 2009; Girju et al., 2006).
Recent advances incorporate LLMs prompting to enhance taxonomy construction. For example, Wan et al. (2024); Zeng et al. (2024a); Chen et al. (2023); Zeng et al. (2024b) apply zero-/few-shot reasoning and ensemble ranking, while others explore open-ended, vocabulary-free taxonomy creation (Gunn et al., 2024), self-supervised expansion in low-resource domains (Mishra et al., 2024), and graph-based methods leveraging metadata and citations (Cong et al., 2024; Sas and Capiluppi, 2024; Shen et al., 2024). Optimization and in-context learning have also shown promise (Hu et al., 2024b; Shi et al., 2024; Xu et al., 2025; Jain and Espinosa Anke, 2022; Chen et al., 2021).
Our work differs in scope, scale, and methodological design. We focus on scaling taxonomy induction for the domain of scholarly literature—a setting that presents greater challenges than typical setups (e.g., entity hierarchy) due to the complexity, size, and evolving nature of scientific content. The most comparable effort is by Oarga et al. (2024), though our broader objectives require fundamentally different algorithmic strategies and operate without access to ground truth labels.
@@
Structured representation of science: As science grows at an unprecedented rate (Teufel et al., 1999; Pertsas and Constantopoulos, 2017; Constantin et al., 2016; Fisas et al., 2016; Liakata et al., 2010), numerous frameworks have emerged to structure this information through knowledge graphs and taxonomies (Fathalla et al., 2017; Jaradeh et al., 2019; Oelen et al., 2020; Vogt et al., 2020; Soldatova and King, 2006). Recent work includes prompt-based topic modeling (Pham et al., 2024), iterative taxonomy construction that incorporates object properties and graph mining (Cui et al., 2024; Marchenko and Dvoichenkov, 2024), and hybrid approaches that combine curated ontologies with data-driven maps (Zimmermann et al., 2024). Our work builds on these efforts by constructing a high-quality hierarchical structure tailored to scientific literature, in three key ways. The prior work: (1) Produces shallow hierarchies, typically only one or two levels deep; (2) Uses cluster labels based on keywords, whereas ours are derived from natural-language summaries of paper content; (3) Depends heavily on manual effort, while our pipeline is fully automated.
Finally, in Table 1 we summarize the differences with existing hierarchical resources. While most prior systems are limited to one or two levels of depth and rely on manually assigned keyword labels for indexing—a process often prone to bias (Hadfield, 2020)—our approach supports deeper, algorithmically generated hierarchies with semantically rich node descriptions. This enables a more flexible and interpretable representation of scientific knowledge.
@@ 3 Science Hierarchography: Toward Hierarchy of Scholarly Work
We begin with a formal problem definition (§3.1), followed by content representation (§3.2) and depth considerations (§3.3).
@@ 3.1 Formal Problem Statement
Formally, this goal is defined as inferring a graph that encodes relationships between nodes (atomic concepts representing scholarly ideas or goals). This is formally defined as $T\vcentcolon=(N,E)$ where:
  • NN is the set of all concepts (nodes) present in the hierarchy: $N=\{n_{1},n_{2},\ldots,n_{N}\}$.
@@
  • For each edge $e\in E$, there is a type attribute: $\text{type}(e)\in\{\text{``isA''},\text{``instanceOf''}\}$ associated with the edge.
@@
The edge (relations connecting two nodes) have types that is either “isA” relationships or “instanceOf.” The "isA" relationship defines a hierarchical link between node pairs, indicating a child node is a subclass of its more abstract parent node (e.g., "RLHF isA RL" means "RLHF" is a type of "RL"). The "instanceOf" relationship indicates that a specific instance belongs to a concept category (e.g., "Ouyang et al. (2022) instanceOf RLHF" means "Ouyang et al. (2022)" is a specific instance of the broader “RLHF” cluster).
3.2 Decomposing Papers to Contributions
A central challenge is how to represent the content of scholarly work within hierarchy nodes. Scientific papers are idea-dense, often combining broad goals, specific problems, and technical methods. To capture this complexity, we extract structured representations that disentangle these distinct aspects (D’Souza and Auer, 2020). This also mitigates the issue of input length: papers typically range from 4 to 10 pages (5K to 10K tokens), making full-document processing across large corpora infeasible and costly for LLMs.
@@ 3.2 Decomposing Papers to Contributions
We use an LLM to preprocess each paper (title and abstract) and break them down into a pre-defined set of contributions, akin to prior work Hope et al. (2017); Chan et al. (2018) that mine “problem schema” from existing documents. We consider the following contribution types: (1) problem statement (the problem addressed), (2) solution (the technical approach used), (3) result (the key finding), and (4) topic (the overarching themes). (See §C for prompts and examples). We note that each contribution may include additional dimensions (sub-contributions). For instance, a “result” encompasses both the “outcome” and its “potential impact.”
3.3 Choosing Hierarchy Depth
While the ideal number of hierarchy layers is ultimately empirical, we can build useful intuition from the structure of a near-balanced tree. For a tree with branching factor $b$ and depth $L$, the total number of nodes is roughly $O(b^{L})$. To organize $C$ contributions, the number of nodes should scale with $C$, implying a depth of $L=O(\log_{b}C)$. In practice, we use $L=3$ for a 2K-paper corpus and $L=4$ for 10K papers, consistent with this logarithmic scaling. Extrapolating further, corpora of $10^{7}$ papers would likely require depths of $L=6$ or $7$.
4 Tackling Science Hierarchography
@@
We present algorithms to address our proposed goal. We start with our main method, Scychic4.1), explore its special cases (§4.2), and then describe alternative baselines that rely more heavily on LLMs (fLMSci; §4.3). While all approaches leverage LLMs to some extent, they differ significantly in their reliance on them: some require many calls (linear or quadratic in the number of papers), while others are more efficient (e.g., logarithmic). Since our goal is to scale to millions of papers, minimizing LLM usage is critical. Our objective is to identify the method that yields the highest-quality hierarchy with the lowest LLM overhead, balancing quality, latency, and cost.
@@ 4.1 Scychic: Alternating Between Clustering and Summarization
For the first layer ($L=1$), we partition the entire paper collection into $k_{1}$ top-level clusters using their embeddings. For the second level, we process each top-level cluster independently, i.e., within each top-level cluster, we take only the papers belonging to that cluster and apply clusterer to their embeddings to divide them into subgroups.
We repeat this process for all layers from $L$ to $\left\lfloor L/2\right\rfloor+1$. To build layer $l$, we start by embedding the generated cluster summaries from the level below l1l-1 using embedder, similar to how we embedded the papers. We then run the clustering clusterer on these new embeddings and generate abstracted summaries for the clusters to group these summaries into higher-level clusters. This bottom-up aggregation continues until we connect with the previously constructed level $\left\lfloor L/2\right\rfloor$ clusters.
@@
7:   for each cluster from layer l1l-1 do